OpenStudy (anonymous):

Simplify /compute sin(0.7t) + cos(1.1t) +sin(1.5t) + cos(1.9t) + sin(2.3t) + cos(2.7t) + sin(3.1t) + cos(3.5t) +sin(3.9t) =

OpenStudy (anonymous):

try grouping like this sin(0.7t) + sin(2.3t) + cos(1.1t) + cos(1.9t) + sin(1.5t) + cos(3.5t) + sin(3.1t) + sin(3.9t) + cos(2.7t) :P

OpenStudy (anonymous):

Yes, go on please.

OpenStudy (anonymous):

ah.. leave it :P

OpenStudy (anonymous):

It will be solved eventually.

OpenStudy (anonymous):

it maybe solved :P not sure by my side

OpenStudy (anonymous):

I give you hope. (and change ?...)

OpenStudy (anonymous):

@mathslover welcome help

OpenStudy (anonymous):

Well, no change so far

OpenStudy (anonymous):

Hey there @mathslover you multiplied (and grown in NUMBER !) congrats

OpenStudy (anonymous):

Any ideas people ?

OpenStudy (anonymous):

I wouldn't ask this question unless it was educational and elegant

OpenStudy (anonymous):

No luck bro :P

OpenStudy (anonymous):

Well what about this idea @vikrantg4 @mathslover \[\cos(\alpha) = 0.5(e^{i \alpha} + e^{-i \alpha})\]

OpenStudy (anonymous):

omg i don't know details in complex numbers lool

OpenStudy (anonymous):

\[\sin(\beta) = -i 0.5(e^{i\beta} - e^{-i\beta})\]

OpenStudy (anonymous):

Forget complex numbers in general use this formula only

OpenStudy (anonymous):

it's gonna be a pain to use this in 9 expresions

OpenStudy (anonymous):

The whole point is IT IS NOT PAIN there is a way

OpenStudy (anonymous):

lool :PP

OpenStudy (anonymous):

@zzr0ck3r I gave a guidance to solution

OpenStudy (anonymous):

@vikrantg4 Try using your own first advice and these

OpenStudy (zzr0ck3r):

excuse me?

OpenStudy (anonymous):

I thought you might help @zzr0ck3r

OpenStudy (anonymous):

Ok time to spoil or not yet ?

OpenStudy (anonymous):

It' getting more complex using my method

OpenStudy (anonymous):

Replace everything with exponents - then what do you get ?

OpenStudy (anonymous):

Lets say\[e^{i\alpha} + e^{i 2\alpha} = ?\]

OpenStudy (anonymous):

Hi @ajprincess

OpenStudy (anonymous):

Thanks Flute-passionate @mathslover

mathslover (mathslover):

.welcome. ;)

OpenStudy (anonymous):

Well ladies and gentleman - what is this\[e^\sigma + e^{2\sigma} + e^{3\sigma} + e^{4\sigma}\]

OpenStudy (anonymous):

It is a .....it is Ahhh ....

mathslover (mathslover):

expression. For sure..

OpenStudy (anonymous):

geometric sequence

OpenStudy (anonymous):

;\

OpenStudy (anonymous):

Great @Coolsector , NOW ?

OpenStudy (anonymous):

q=e^a

OpenStudy (anonymous):

Yes but what do we DO now ?

OpenStudy (anonymous):

using the formula for sum

OpenStudy (anonymous):

Step forward - what will it be in this case ?

OpenStudy (anonymous):

thx @Coolsector

OpenStudy (anonymous):

well this is great.. thank you for posting this

OpenStudy (anonymous):

how can i do it ? :| lol i have no idea

OpenStudy (anonymous):

http://www.wolframalpha.com/

OpenStudy (anonymous):

Hi @sauravshakya i must admit, pleased to see you here

OpenStudy (anonymous):

Thanx and GOOD ONE!

OpenStudy (anonymous):

In fact there are A LOT of practical applications of this 1) Z-transform 2) Digital Signal Processing filters design 3) Complex Integration . . .

OpenStudy (anonymous):

Thanks @TheViper pal !

OpenStudy (theviper):

\(\Huge{\color{violet}{\ddot{\smile}}}\)

OpenStudy (anonymous):

You tell me in PM how one does THAT !

OpenStudy (theviper):

PM ?

OpenStudy (anonymous):

Private messg.

OpenStudy (anonymous):

\[ \Huge{\color{green}{\ddot{\smile}}} \]

OpenStudy (anonymous):

Mikael 4 Finding the sum of a trigonometric series uses De Moivre's Theorem, geometric progression, trigonometric identities, and complex numbers. Top universities for mathematics test prospective students by setting questions in their entrance tests that stretch the most capable students. The example used here has been used in the past by Cambridge University and other top math colleges. The question is an excellent one, that uses De Moivre's Theorem, geometric progressions, trigonometric identities, complex numbers, and algebraic manipulation. Problem: Find the General Solution to cos(x) + cos(2x) + cos(3x) + ... + cos(nx) The core of the solution is to recognise that cos.x + cos.2x + cos.3x + ... + cos.nx = Real part of (cos.x + i.sin.x + cos.2x + i.sin.2x + cos.3x + i.sin.3x + ... + cos.nx + i.sin.nx) (where i = √-1) = Real part of [ (cos.x + i.sin.x) + (cos.x + i.sin.x)^2 + (cos.x + i.sin.x)^3 + ... + (cos.x + i.sin.x)^n ] ... [Eq'n 1] This is a geometric progression: let Y = (cos.x + i.sin.x), and the series becomes Real part of (Y + Y^2 + Y^3 + ... + Y^n) From the rules of a geometric progression, this is = Real part of [Y.(Y^n - 1) / (Y - 1)] Now substitute Y = (cos.x + i.sin.x) back in again, and the sum becomes = Real part of (cos.x + i.sin.x).[ (cos.x + i.sin.x)^n –1] / (cos.x + i.sin.x - 1) = Real part of (cos.x + i.sin.x).[ (cos.nx + i.sin.nx) –1] / (cos.x + i.sin.x - 1) Rationalise The Denominator The denominator is complex. To get rid of the “i” term, multiply numerator and denominator by the complex conjugate of the denominator, (cos.x –1 – i.sin.x), and the summation becomes =(cos.x + i.sin.x).[ (cos.nx + i.sin.nx) –1].(cos.x -1 + i.sin.x ) / [ (cos.x -1 + i.sin.x ).( cos.x -1 - i.sin.x) ] For clarity, the denominator will be expanded first: = (cos.x -1 + i.sin.x ).( cos.x -1 - i.sin.x) This is simply the difference of two squares: = (cos.x – 1)^2 – (i.sin.x)^2 = (cos.x)^2 – 2.cos.x. + 1 – (i^2.(sin.x)^2) = (cos.x)^2 – 2.cos.x. + 1 + (sin.x)^2 But (cos.x)^2 + (sin.x)^2 = 1, so the denominator becomes = 2 – 2.cos.x =2.(1 – cos.x) Evaluate the Numerator The numerator is = (cos.x + i.sin.x).[ (cos.nx + i.sin.nx) –1].( cos.x -1 - i.sin.x) = (cos.x.cos.nx + cos.x.i.sin.nx - cos.x + i.sin.x.cos.nx + i.sin.x.i.sin.nx - i.sin.x)).( cos.x -1 - i.sin.x ) But from the standard trigonometric identities, (cos.x.cos.nx - sin.x.sin.nx) = cos(nx + x) = cos(n+1).x and cos.x.i.sin.nx + i.sin.x.cos.nx = sin(nx + x) = sin(n+1).x So the numerator becomes = (cos(n+1).x + i.sin(n+1).x - cos.x - i.sin.x). ( cos.x - 1 - i.sin.x ) = cos.x.(cos(n+1).x + i.sin(n+1).x - cos.x - i.sin.x) - (cos(n+1).x + i.sin(n+1).x - cos.x - i.sin.x) - i.sin.x.(cos(n+1).x + i.sin(n+1).x - cos.x - i.sin.x) = cos.x.cos(n+1).x + i.cos.x.sin(n+1).x - cos².x - i.cos.x.sin.x - cos(n+1).x - i.sin(n+1).x + cos.x + i.sin.x - i.sin.x.cos(n+1).x - i².sin.x.sin(n+1).x + i.sin.x.cos.x + i².sin.x.sin.x = cos.x.cos(n+1).x + i.cos.x.sin(n+1).x - cos².x - i.cos.x.sin.x - cos(n+1).x - i.sin(n+1).x + cos.x + i.sin.x - i.sin.x.cos(n+1).x + sin.x.sin(n+1).x + i.sin.x.cos.x - sin².x ... [Eq'n 2] = cos.x + cos.2x + ... cos.nx + i.(sin.x + sin.2x + ... + sin.nx) Find cos.x + cos.2x + cos.3x + ... + cos.nx From Equation [1], the sum of the series is the real part of the geometric progression: Gathering real terms, the real part of equation 2 = cos.x.cos(n+1).x - cos².x - cos(n+1).x + cos.x + sin.x.sin(n+1).x - sin².x Simplifying, cos.x.cos(n+1).x + sin.x.sin(n+1).x = cos[(n+1).x - x] = cos.nx, and -cos²x - sin²x = -(cos²x + sin²x) = -1, so the real part becomes = cos.nx - cos.(n+1).x - 1 + cos.x So the sum of cos.x + cos.2x + ... + cos.nx = numerator / denominator = (cos.nx - cos.(n+1).x - 1 + cos.x) / 2.(1 - cos.x) = (cos.nx - cos.(n+1).x) / (2 - 2.cos.x) - 1 / 2 Q.E.D. To verify this, choose x = π / 5 and n = 3 as an example: (cos.3.π / 5 - cos.(4.π / 5)) / (2 - 2.cos.π / 5) - 1 / 2 = 0.809 cos(π / 5) + cos(2.π / 5) + cos(3.π / 5) = 0.809 + 0.309 - 0.309 = 0.809 Sum of Trigonometric Series Using De Moivres Theorem: cosx+cos2x+ | Suite101.com http://suite101.com/article/sum-of-trigonometric-series-using-de-moivres-theorem-cosxcos2x-a369345#ixzz27BKBaewU

OpenStudy (anonymous):

cos(3.5t) +sin(3.9t) cos(3.9t-0.4t) +sin(3.9t) cos(3.9t) cos(0.4t) +sin(3.9t) sin(0.4t) +sin(3.9t) Like this you can reduce all the terms to sin/cos(3.9t) multiplied by a constant.

OpenStudy (anonymous):

Well @henpen you are correct (in your general hunch) but guess what - the most elegant and short way to prove and UNDERSTAND addition formulas is BY De Moivre theorem !

OpenStudy (anonymous):

I do understand that, just was too lazy to read the full proof.

OpenStudy (anonymous):

Oddly enough I encountered de Moivre's theorem before any addition formulae. Anyway, thanks for bringing me here again- the proof is very elegant!

OpenStudy (anonymous):

Hey You sooth my tormented soul. This place is very double edged regarding how people react to real elegant things shown them.... For example this stinking sid whatever http://openstudy.com/study#/updates/505e1877e4b02e1394101626